A sample of 200 claims processed under the the insurance system revealed 180 of the claims were paid in full. Is there evidence, at the .05 level of significance that the population proportion of claims processed under the new system is higher than the article reported the previous system and then compute the p-value and interpret its meaning.








Answer by stanbon(75887)

A sample of 200 claims processed under the the insurance system revealed 180 of the claims were paid in full. Is there evidence, at the .05 level of significance that the population proportion of claims processed under the new system is higher than the article reported the previous system and then compute the p-value and interpret its meaning.
------------------------
Ho: p <= ?
Ha: p > ?
---
p-hat = 180/200 = 0.9
---
z(0.9) = (0.9-?)/sqrt(?*(1-?)/200) = ?????
etc.
---
The previous reported proportion is needed in your post.
Cheers,
Stan H.